OpenStudy (anonymous):

A basketball was thrown from one end of a court to another. It had an initial upward velocity of 48 feet per second. How long did it take for the ball to hit the ground? Formula (h=v-16t^) a.) 3 Seconds b.)0.3 Seconds c.)-3 Seconds d.)-0.3 Seconds

OpenStudy (nurali):

v=velocity=48, h=height=0, t=time h = vt – 16t^2 48t-16t^2=0 16t(3-t)=0 -t=-3 t=3 sec

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (nurali):

Anytime.